Steam from a superheater at 7 bar, 300 degree Celsius is mixed in steady adiabatic flow with wet steam at 7 bar, dryness fraction 0.9.calculate the mass of wet steam required per kilogram of superheated steam to produce steam at 7 bar, dry saturated.
1
Expert's answer
2021-09-14T01:53:53-0400


We have x amount of superheated steam at 7 bar ("700kpa" ) and "300C=> h=3059kj\/kg" (h of superheated at temp and press)


total energy of the closed system:

"x\times3059+y\times2556.22=z\times2762.8; z=x+y" (closed system)

"x\times3059+y\times 2556.22=(x+y)\times2762.8" ; lets assume L is the mass ratioof wet steam to superheated

=> "L=y\/x" "x(3059+L(2556.22)=x(1+L)2762.8"

"(3059+L\times2556.22)=(1+L)\times2762.8"

solve for L and you should get "1.43 kg" wet steam/kg superheated (itwill be a little off because my steam tables might be different orrounding)